\begin{thebibliography}{99}
 
\bibitem{AMSW} Adler, R.J., Monrad, D., Scissors, R.H. \& Wilson, R. (1983): Representations, decompositions and sample
function continuity of random fields with independent increments. Stoch. Process. Appl. 15(1),3-30.
  
\bibitem{CMV} Capelas de Oliveira, E.; Mainardi \& F.Vaz, J. (2011): Models based on Mittag Leffler functions for anomalous relaxation in dielectrics. ariXir: 1106.1761  v2[cond-mat.Stat-mech] 13 Feb. 2014.

\bibitem{DW} Dalang, R. C. \& Walsh, J. B. (1992): The sharp Markov property of L\'evy sheets. The Annals of Probability 20 (2), 591-626.

\bibitem{DOP1} Di Nunno, G., \O ksendal, B. \& Proske, F. (2004):
White noise analysis for L\' evy processes. 
J. Functional Analysis 206 (2004),109 -148.

  \bibitem{DOP} Di Nunno, G., \O ksendal, B. \& Proske, F. (2009): Malliavin Calculus for L\' evy Processes with Applications to Finance. Springer.
  
 \bibitem{DMO} Draouil, O., Moulay-Hachemi, Y. \& \O ksendal, B. (2025): The stochastic heat inclusion with fractional time driven by  time-space Brownian noise. Submitted.
 	arXiv:2506.08275
  \bibitem{fedotov2007}
Fedotov, S. and Iomin, A. (2007).
  Migration and proliferation dichotomy in tumor cell invasion.
 Physical Review Letters, 98(11):118101.  
    
 \bibitem{GV} Gel'fand, I. M. \& Vilenkin, N. Ya.(1971): Generalized Functions. Vol. 4. Academic Press [Harcourt Brace Jovanovich Publishers], New York.
 
\bibitem{H} Hida, T. (1980): Brownian Motion. Springer, New York.

\bibitem{HKPS} Hida, T., Kuo, H-H., Potthoff, J. \& Streit, L. (1993): White Noise: An Infinite Dimensional Calculus. Springer, New York.

   \bibitem{Hu} Hu, Y. (2019): Some recent progress on stochastic heat equations. Acta Mathematica Scientia 39B(3); 874-914.
   
\bibitem{HOUZ}
Holden, H., \O ksendal,B., Ub\o e, J. \& Zhang, T. (2009)  Stochastic Partial Differential Equations, Springer.

 \bibitem{IR} Iafrate, F. \& Ricciutu, C.(2024): Some families of random fields related to multiparameter L\' evy processes. J. Theoretical Probability 37: 3055-3088. https://doi.org/10.1007/s10959-024-01351-3.
 
 \bibitem{Kilbas} Kilbas, A.A., Srivastava,H.M. \& Trujillo, J. J. (2006): 
	 {\emph{ Theory and Applications of Fractional Differential Equations }}, Elsevier  Science B.V, Amsterdam. 
     
\bibitem{KKS} Kochubel, A. N., Kondratiev, Y. \& da Silva, J. L. (2021): On fractional heat equation. Fractional Calculus \& Applied Analysis 24 (1), 73-87.
	     
   \bibitem{MS} Meerschaert, Mark M., Sikorskii, Alla (2019): Stochastic Models for Fractional Calculus, $2^{nd}$ edition. De Greuter.
   
   \bibitem{ML} Mittag-Leffler, M.G. (1903): Sur la nouvelle fonction E(x). \emph{C. R. Acad. Sci. Paris} {\bf 137} 554--558.
   
  \bibitem{MO1} Moulay Hachemi, R.Y.  \& \O ksendal, B. (2023): The fractional stochastic heat equation
driven by time-space white noise. Fractional Calculus \& Applied Analysis 26, 513-532. Springer.
https://doi.org/10.1007/s13540-023-00134-7.

\bibitem{murray2003}
Murray, J.D. (2003).
 Mathematical Biology II: Spatial Models and Biomedical Applications (3rd ed.). Springer-Verlag.
\bibitem{NualartSchoutens} Nualart, D. \& Schoutens, W. (2000): Chaotic and predictable representations for L\'evy processes.
Stochastic Process. Appl., 90(1):109-122.


\bibitem{OP} \O ksendal, B. \& Proske, F. (2004):
White noise of Poisson random measures. 
Potential Analysis 21 (2004), 375-403.

\bibitem{PK} Patel, A. \& Kosko, B. (2007): Levy Noise Benefits in Neural Signal Detection, 2007 IEEE International Conference on Acoustics, Speech and Signal Processing - ICASSP '07, Honolulu, HI, USA, 2007, pp. III-1413-III-1416, doi: 10.1109/ICASSP.2007.367111.  

\end{thebibliography}

\documentclass[12pt]{article}%
\usepackage[applemac]{inputenc}
\usepackage{graphicx}  
\usepackage{caption}   
\usepackage{amsmath,amssymb}
\usepackage{amsfonts}
\usepackage[applemac]{inputenc}
\usepackage{amsmath,amssymb}
\usepackage{color}
\usepackage{color, colortbl}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{tikz}
\usepackage{geometry}
\usetikzlibrary{arrows}
\usepackage{amsfonts}
\usepackage{amsmath,amssymb}
\usepackage[applemac]{inputenc}
\usepackage{color}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{tikz}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newcommand{\E}{E}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{definition}[theorem]{Definition}

\newtheorem{example}[theorem]{Example}

\newtheorem{remark}[theorem]{Remark}

\usetikzlibrary{arrows,shapes,automata,petri}
\newenvironment{myenumerate}{

\begin{enumerate}}{\end{enumerate}}
\newcommand{\dproof}{\noindent {Proof.} \quad}
\newcommand{\fproof}{\hfill $\square$ \bigskip}

\numberwithin{equation}{section}

\definecolor{LightCyan}{rgb}{0.88,1,1}
\def\cF{{\mathcal {F}}}

\def\RR{{\mathbb{ R}}}
\def\cB{{\mathcal{B}}}

\def\1B{\text{1\!\!I}}

\def\l{\langle}
\def\<{\langle}
\def\>{\rangle}
\def\P{\mathbb{P}}
\def\R{\mathbb{R}}
\def\l{\lambda}
\def\S{\mathcal{S}}
\def\E{\mathbb{E}}
\def\N{\mathbb{N}}

\def\F{\mathcal{F}}

\def\realio{\mathbb{R}_{0}}

\begin{document}

\title{Multiparameter L\'evy white noise theory and applications}

\author{Olfa Draouil$^{1}$, Rahma Yasmina Moulay Hachemi$^{2}$ \& Bernt \O ksendal$^{3}$
}
\date{29 October 2025}
\maketitle 
\footnotetext[1]{Department of Mathematics, University of Tunis El Manar, Tunisia\newline
Email: olfa.draouil@fst.utm.tn}

\footnotetext[2]{%
Department of Mathematics, University of Saida, Algeria.\newline
Email: yasmin.moulayhachemi@yahoo.com}

\footnotetext[3]{%
Department of Mathematics, University of Oslo, Norway. \\
Email: oksendal@math.uio.no.}
\paragraph{MSC [2020]:}
\emph{30B50; 34A08; 35D30; 35D35; 35K05; 35R11; 60H15; 60H40}

\paragraph{Keywords:}
\emph{L\' evy sheets, compensated Poisson random measures, white noise calculus.}

\begin{abstract}
We construct a white noise theory and white noise calculus for the (multi-parameter) L\' evy sheet and its compensated Poisson random measures. The theory applies to stochastic partial differential equations subject to L\' evy noise.
  \end{abstract}

\section{Introduction}

The theory of white noise for Brownian motion was initiated by T.Hida \cite{H} and further developed by him and coauthors Kuo, Potthoff and Streit in \cite{HKPS}. Subsequently an extension to (multiparameter)  Brownian sheet was carried out and applied to stochastic partial differential equations by Holden, Ub\o e, \O ksendal and Zhang  in \cite {HOUZ}. 

Later Di Nunno, \O ksendal \& Proske \cite{DOP1} and \O ksendal \& Proske \cite{OP} developed a white noise theory for L\' evy processes and for compensated Poisson random measures, respectively. See the book \cite{DOP} for a comprehensive coverage of this.  

In view of the success of this white noise theory for the Brownian sheet and the L\' evy process it is natural to ask if it is also possible to carry out a construction of white noise for the \emph{(multiparameter) L\' evy sheet}. 

The theory of L\' evy sheets was developed by Iafrate \& Ricciutu  in \cite{IR} and by Dalang \& Walsh \cite{DW}. See also the earlier paper Adler et al  \cite{AMSW}. 
But, to the best of our knowledge, so far there has been no papers dealing with the white noise theory for L\' evy sheets. The purpose of this paper is to provide such a construction and give some applications.  This theory appears to be new. We obtain it by combining the results in the papers \cite{IR}, \cite{DW}, \cite{DOP} and \cite{OP} above.



\section{Multiparameter L\'{e}vy white noise and stochastic distributions}
This is the main section, in which we first give a construction of multiparameter L\'{e}vy sheets, and then the associated white noise theory and calculus.

First we recall the theory of L\' evy sheets, mainly based on \cite{IR}, \cite{DW}:

\subsection{Multiparameter L\'{e}vy sheets}

\begin{definition}
  A \emph{multiparameter L\'{e}vy sheet} is a stochastic process $L(x)=L(x,\omega); x \in \R_+^n,\omega \in \Omega$ defined on a probability space $(\Omega,\F,\P)$ and satisfying the following conditions:
\begin{itemize}
\item
    $L$(0)=0 a.s.
\item
$L$ has stationary, independent increments with respect to the natural partial ordering on $\R_+^n$
\item
$L$ is continuous in probability and has c\'adl\'ag paths a.s. 
\end{itemize}
\end{definition}

\begin{remark}
If $R=\Pi_{j=1}^n (x_j,x'_j]$ is a box in $\R_+^n$ then $\Delta_R L$ denotes the increment of $L$ over the box $R$. For example, if $n=2$ then
$$\Delta_R L= L(x_1',x_2')-L(x_1',x_2) - L(x_1, x_2') + L(x_1,x_2).$$
\end{remark}
\begin{remark}    
If we also assume that $L$ has continuous paths, then $L$ becomes a \emph{Brownian sheet}.
\end{remark}

\begin{definition}
\begin{myenumerate}
    \item
Let $R$ denote a box in $\R_+^n$ and let $U$ be a Borel subset of $\R_0:=\R \setminus 0$.
Define \emph{the jump measure} $N(R,U)$ to be the number of jumps of $L$ within the box $R$ of jumps size in $U$. By stationarity the map $R \mapsto N(R,U);  U \in \mathcal{B}(\R_0)$ depends only on the volume of $R$ and we can write the random measure generated by $(R,U) \mapsto N(R,U)$ in differential form as  $N(dx, du)$, where $dx$ and $du$ are the Lebesgue measure differentials corresponding to $R$ and $U$, respectively.
\item
Put $R_1=[0,1]^n$. The measure $U \mapsto \nu(U):=\E [N(R_1,U)];  U \in \mathcal{B}(\R_0)$ is called the L\'{e}vy measure of $L$.
\item
The corresponding \emph{compensated jump measure} is defined by
$$ \widetilde{N}(dx,du)= N(dx,du) - \nu(du) dx,$$
where $dx=dx_1 dx_2 ... dx_n$ is the volume differential at $x=(x_1, x_2, ... ,x_n) \in \R^n$.
\end{myenumerate}
\end{definition}

In the following we will for simplicity assume that
\begin{align}
\int_{\R_0} z^2 \nu(dz) < \infty.
\end{align}
Then by the L\'{e}vy-Khinchine theorem the L\'{e}vy sheet $L$ can be given the following representation:
\begin{align}
    L(x)=\alpha \Pi_{j=1}^n x_j + \sigma B(x) + \int_{\R_0} z \widetilde{N}(x, dz); \ x \in \R_+^n,
\end{align}
for some constants $\alpha, \sigma$, where $B(x)$ is a Brownian sheet.

In particular, if we assume that $\alpha=\sigma=0$ then we have the \emph{pure jump L\'evy sheet}, with representation
\begin{align} 
L(x)=\int_{\R_0} z \widetilde{N}(x,dz)=\int_0^x \int_{\R_0} \widetilde{N}(d\xi,dz); \ x \in \R_+^n. \label{LN}
\end{align}

\subsection{The L\'{e}vy white noise probability space and the pure jump L\'{e}vy sheet}

In the following we let $\S=\S(\R^n)$ denote the Schwartz space of rapidly decreasing smooth functions on $\R^n$, and we let $\S'=\S'(\R^n))$ denote its dual, which is the space of tempered distributions on $\R^n$, equipped with the weak$^{*}$ topology. 

It is possible to construct a \emph{pure jump L\'{e}vy sheet} $L$ corresponding to a given  L\'{e}vy measure $\nu$  
as a process on the Borel $\sigma$-algebra $\cF = \cB(\S'(\R^n))$. The details are the following:

Let $\nu$ be a measure on $\R_0$ such that
\begin{align}
\int_{\R_0} z^2 \nu(dz) < \infty.
\end{align}
By the Bochner-Minlos-Sazonov theorem\index{Bochner-Minlos-Sazonov theorem} (see \cite{GV}), there exists a probability measure $\P$ on the Borel $\sigma$-algebra $\cF = \cB(\S'(\R^n))$ such that
\begin{equation}
\int_\Omega e^{i\langle\omega,\varphi \rangle}\P(d\omega)=\exp\left(\int_{\R^n}\Psi(\varphi(x))dx\right);
\quad \varphi\in\mathcal{S}(\R^n), \label{boch}
\end{equation}
where $i=\sqrt{-1}$ and $\langle\omega,\varphi\rangle$ denotes the action of $\omega\in \Omega:=\mathcal{S}'(\R^n)$ on $\varphi\in \mathcal{S}(\R^n)$ and 
\begin{equation}
\Psi(w)=\int_{\R}\left(e^{i\,w\,z}-1-i\,w\,z\right)\nu(dz); w \in \R.
\end{equation}

We let $\E[\cdot]=\E_{\P}[\cdot]$ and $Var[\cdot]=Var_{\P}[\cdot]$ denote expectation and variance, respectively, with respect to $\P$.\\
Note that by \eqref{boch} we have, for all parameters $a \in \R$,
\begin{align}
    \E| e^{i\langle \omega,a \varphi \rangle)}]=\exp\left(\int_{\R^n}\int_{\R_0} \{ e^{i a \varphi(x)z} - 1 - i a \varphi(x)z\} \nu(dz) dx \right) \label{boch2}
\end{align}
Expanding the exponential function on both side of \eqref{boch2} and comparing the terms with the same power of $a$, we obtain the following result:
\begin{lemma}
\label{2.5}
Let $\varphi\in\mathcal{S}(\R^n)$. 
Then we have
\begin{equation}
\E[\langle\cdot,\varphi\rangle]=0
\end{equation}
and
\begin{equation}
Var[\langle\cdot,\varphi \rangle]:=\E[\langle\cdot,\varphi\rangle^2] = M \int_{\R^n} \varphi^2(x)dx, \label{Var}
\end{equation}
where
\begin{align}
    M=\int_{\R_0} z^2 \nu(dz).
\end{align}
Therefore,
\begin{align}
    \E[(\langle \omega,\varphi_1\rangle - \langle\omega,\varphi_2\rangle)^2]=M \int_{\R^n} (\varphi_1(x)-\varphi_2(x))^2 dx=M \| \varphi_1 - \varphi_2 \|^2_{L^2(\lambda)}, 
\end{align}
where $\lambda$ is Lebesgue measure on $\R^n$.

\end{lemma}

Using this result we can for each  $\omega \in \Omega$ extend the definition of $\langle \omega,f\rangle$ from $f\in\mathcal{S}(\R^n)$ to any $f\in L^2(\lambda)$ as follows:

If $f\in L^2(\lambda)$, choose $f_n\in\mathcal{S}(\R^n)$ such that $f_n\rightarrow f$ in $L^2(\lambda)$. 
Then by Lemma \ref{2.5} we see that $\{\langle
\omega,f_n\rangle\}_{n=1}^\infty$ is a Cauchy sequence in $L^2(\P)$ and hence convergent in $L^2(\P)$. 
Moreover, the limit depends only on
$f$ and not the sequence $\{f_n\}_{n=1}^\infty$. 
We denote this limit by $\langle \omega,f\rangle; \  \omega \in \Omega, f \in L^2(\lambda).$

In particular, if we define
$$\widetilde{L}(x):=\widetilde{L}(x,\omega):=\langle\omega,\chi_
{\Pi_{j=1}^n [0,x_j]} \rangle;\quad x=(x_1, x_2, ... ,x_n) \in \R_+^n,$$
then we have the following result:

\begin{theorem}
The stochastic process $\widetilde{L}(x)$, $x \in \R_+^n$ has a c\`{a}dl\`{a}g version, denoted by $L$. 
This process $L(x)\ ; x\in \R_+^n$ is a (pure jump) L\'{e}vy
sheet with L\'{e}vy measure $\nu$.
\end{theorem}

\subsection{A chaos expansion}

We now proceed to define the white noise of the L\'{e}vy sheet we constructed in the previous section. We will follow the presentation in \cite{DOP} Section 13.2  closely, but with the necessary modifications due to the fact that we are now dealing with an $n$-parameter L\'{e}vy sheet, not a classical 1-parameter L\'{e}vy process. Here are the details:\\

From now on we assume that the L\'{e}vy measure $\nu $ satisfies the
following condition:\\
For all $\varepsilon >0$ there exists $\lambda >0$ such that
\begin{equation}
\int_{\realio\backslash (-\varepsilon ,\varepsilon )}\exp (\lambda
\left\vert z\right\vert )\nu (dz)<\infty   \label{nu}
\end{equation}
One can prove that if this holds, then $\nu $ has finite moments of all orders $n\geq 2$. 
For example, the condition clearly holds if $\nu $ is supported on $[-C,C]$ for some $C>0$.

From this condition one can prove that the polynomials are dense in $L^{2}(\rho )$,
where
\begin{equation}
\rho (dz)=z^{2}\nu (dz).  
\end{equation}
See \cite{NualartSchoutens}. We let $\left\{\eta_{m}\right\} _{m\geq 0}=\left\{ \eta_0 =1,\eta_{1},\eta_{2},...\right\} $ \label{simb-1303} be the orthogonolization of
$\left\{1,z,z^{2},...\right\} $ with respect to the inner product of $L^{2}(\rho )$.

Define
\begin{equation}
p_{j}(z):=\left\Vert \eta_{j-1}\right\Vert _{L^{2}(\rho )}^{-1}z\eta_{j-1}(z);
\text{ }j=1,2,...  \label{10.3}
\end{equation}
and
\begin{equation}
m_{2}:=\left( \int_{\realio} z^{2}\nu (dz)\right) ^{\frac{1}{2}}
=\left\Vert \eta_{0}\right\Vert _{L^{2}(\rho )}
=\left\Vert 1\right\Vert_{L^{2}(\rho )}.  \label{10.4}
\end{equation}
In particular,
\begin{equation}
p_{1}(z)=m_{2}^{-1}z\text{ or }z=m_{2}p_{1}(z).  \label{10.5}
\end{equation}
Then $\left\{ p_{j}(z)\right\} _{j\geq 1}$ is an \emph{orthonormal basis}
for $L^{2}(\nu )$.

Recall that the {\it Hermite polynomials\/} $h_n(x); x \in \R$ are defined by
$$h_n(x)=(-1)^n
e^{{1}/{2}x^2}\frac{d^n}{dx^n}(e^{-{1}/{2}x^2});\quad
n=0,1,2,\cdots.\leqno{(2.2.1)}$$
Thus the first Hermite polynomials are
\begin{align}
h_0(x)&=1,\;h_1(x)=x,\;h_2(x)=x^2-1,\;h_3(x)=x^3-3x,\cr
h_4(x)&=x^4-6x^2+3,\;h_5(x)=x^5-10x^3+15x,\cdots.
\end{align}
The \emph{Hermite functions} $\xi_n(x)$ are defined by
$$\xi_n(x)=\pi^{-{1}/{4}}((n-1)!)^{-{1}/{2}}e^{-{1}/{2}x^2}h_{n-1}
(\sqrt{2}x);\quad n=1,2,\cdots.$$
We note the following  useful properties of $\xi_n$:
\begin{itemize}
\item
$\xi_n\in{\cal S}(\R)\quad\text{for all}\;n.$
\item
 The collection $\{\xi_n\}^\infty_{n=1}$ constitutes an
orthonormal basis for $L^2(\R).$
\item
$\sup\limits_{x\in\R}|\xi_n(x)|=\mathcal{O}(n^{-{1}/{12}}).$
\end{itemize}
 Proofs of these statements can be found in Hille and
Phillips (1957), Chapter 21.
\vskip 0.2cm
We will use these functions to define an orthogonal basis for $L^2(\P)$.  In the following, we let
$\beta=(\beta_1,\cdots,\beta_n)$ denote $n$-dimensional multi-indices with
$\beta_1,\cdots,\beta_n\in\N$, where $\mathbb{N}$ denotes the set of all natural numbers.. By the above it follows that the family of tensor
products
$$\xi_\beta:=\xi_{(\beta_1,\cdots,\beta_n)}:=\xi_{\beta_1}\otimes\cdots
\otimes
\xi_{\beta_n}\;;\;\beta\in\N^n$$
forms an orthonormal basis for $L^2(\R^n)$. Let
$\beta^{(j)}=(\beta_1^{(j)},\beta_2^{(j)},\cdots,\beta_n^{(j)})$ be
the $j$th multi-index in some fixed ordering of all $n$-dimensional
multi-indices $\beta=(\beta_1,\cdots,\beta_n)\in\N^n$. We 
may assume that this ordering has the property that
$$i<j\Rightarrow\beta_1^{(i)}+\beta_2^{(i)}+\cdots+\beta_n^{(i)}\leq
\beta_1^{(j)}+\beta_2^{(j)}+\cdots+\beta_n^{(j)},$$
i.e., that the $\{\beta^{(j)}\}^\infty_{j=1}$ occur in increasing order.

Next, define the bijection $\kappa :\mathbb{N}\times \mathbb{N}\longrightarrow
\mathbb{N}$\label{simb-new2} by
\begin{equation}
\kappa (i,j)=j+(i+j-2)(i+j-1)/2;\quad  i,j=1,2, ...
\end{equation}
As in Section 2.2.1 in \cite{HOUZ} we let $\left\{ e _{j}(x)\right\} _{j\geq 1}$ be a fixed ordering of all $n$-times tensor products of the Hermite functions.
This family forms an orthonormal basis of $L^2(\R^n).$
Put
\begin{equation}
\theta _{\kappa (j,k)}(x,z):=e _{j}(x)p_{k}(z); \  x\in \R^n, z \in \R_0.
\end{equation}

In the following we let $\mathcal{J}$ denote the family of all finite multi-indices $\alpha=(\alpha_1, \alpha_2, ..., \alpha_m); \ m=1,2, ...$ of nonnegative integers $\alpha_j$. 

If $\alpha \in \mathcal{J}$ with $Index(\alpha )=j$ and $\left\vert \alpha
\right\vert =m$, we define $\theta ^{\otimes \alpha }: (\R^n \times \R_0)^m \mapsto \R$ by
\begin{eqnarray}
&&\theta ^{\otimes \alpha }(x^{(1)},z_{1};...;x^{(m)},z_{m}) \\
&=&\theta _{1}^{\otimes \alpha _{1}}\otimes ...\otimes \theta _{j}^{\otimes
\alpha _{j}}(x^{(1)},z_{1};... ;x^{(m)},z_{m})  \nonumber \\
&=&\underset{\alpha _{1}\text{ factors}}{\underbrace{\theta
_{1}(x^{(1)},z_{1})\cdot ...\cdot \theta _{1}(x^{(\alpha _{1})},z_{\alpha _{1}})}
}\cdot ...\cdot \underset{\alpha _{j}\text{ factors}}{\underbrace{\theta
_{j}(x^{(m-\alpha _{j}+1)},z_{m-\alpha _{j}+1})\cdot ...\cdot \theta
_{j}(x^{(m)},z_{m})}}.  \nonumber
\end{eqnarray}
We define $\theta _{i}^{\otimes 0}=1.$\\
Finally we let $\theta ^{\hat{\otimes }\alpha }$ denote the
\emph{symmetrized} tensor product\index{tensor product!symmetrized} of the $\theta _{k}$ $^{\prime }s:$
\begin{equation}
\theta ^{\hat{\otimes }\alpha }(x^{(1)},z_{1};...;x^{(m)},z_{m})
=\theta_{1}^{\hat{\otimes }\alpha _{1}}\otimes ...\otimes
\theta _{j}^{\hat{\otimes }\alpha _{j}}(x^{(1)},z_{1};...;x^{(m)},z_{m}).  
\end{equation}
For $\alpha \in \mathcal{J}$ define
\begin{equation}
K_{\alpha }:= I_{\left\vert \alpha \right\vert }\left(
\theta ^{\hat{\otimes }\alpha }\right),
\end{equation}
where in general $I_m$ is the symmetrized iterated integral with respect to $\widetilde{N}$, defined for symmetric functions $g: (\R^n \times \R_0)^m \mapsto \R$ by
\begin{align}
   I_m(g) = \int_{(\R^n \times \R_0)^m } g(x^{(1)},z_1; x^{(2)},z_2; ... ;x^{(m)},z_m) \widetilde{N}^{\otimes m}(dx^{(1)} dz_1; ..., dx^{(m)} dz_m). 
\end{align}

\begin{example}
    For example, if $m=1$ we get
    \begin{align}
        I_1(g)=\int_{\R^n} \int_{\R_0} g(x,z) \widetilde{N}(dx,dz),
    \end{align}
    and if $m=2$ we have
    \begin{align}
        I_2(g)=\int_{\R^n} \int_{\R_0} \Big( \int_{\R^n} \int_{\R_0} g(x^{(1)},z_1; x^{(2)},z_2) \widetilde{N}(dx^{(2)},dz_2) \Big)\widetilde{N}(dx^{(1)},dz_1).
    \end{align}
\end{example}
\begin{example}
\label{Ex10.1}
With $\varepsilon ^{(k)}=(0,...,0,1,0,...)$ with $1$ on $k$ th
place, we have, writing $\varepsilon ^{(\kappa (i,j))}=\varepsilon ^{(i,j)}$,
\begin{equation}
K_{\varepsilon ^{(i,j)}}
= I_1(\theta^{\otimes \varepsilon^{(i,k)}})
=I_{1}\left( \theta_{\kappa (i,j)} \right)
=I_{1}\left( e _{i}(x)p_{j}(z)\right) .  \label{10.11}
\end{equation}%

As in the Brownian sheet case one can prove that $\{ K_\alpha  \}_{\alpha \in \mathcal{J}}$ are
orthogonal in $L^2(\P)$ and
\[
\Vert K_\alpha \Vert^2_{L^2(P)} = \alpha!.
\]
Note that if $|\alpha|=m$,
\[
\alpha ! =\Vert K_\alpha \Vert^2_{L^2(P)} = m! \Vert \theta^{\hat\otimes \alpha} \Vert^2_{L^2((\lambda\times\nu)^m)},
\]
where as before $\lambda$ denotes Lebesgue measure on $\R^n$.
By our construction of $\theta ^{\hat{\otimes }\alpha }$ we know that
any $f\in \widetilde{L}^{2}((\lambda \times \nu )^{m})$ can be written%
\begin{equation}
f(x^{(1)},z_{1};...; x^{(m)},z_{m})=\sum_{\left\vert \alpha \right\vert
=m}c_{\alpha }\theta ^{\hat{\otimes }\alpha
}(x^{(1)},z_{1},..., x^{(m)},z_{m}).
\end{equation}
Hence
\begin{equation}
I_{m}(f_{m})=\sum_{\left\vert \alpha \right\vert =m}c_{\alpha }K_{\alpha }
\end{equation}
\end{example}
This gives the following chaos expansion:
\begin{theorem}
\label{Th10.2}
{\bf (Chaos expansion)}
\newline Any $F\in L^{2}(\P)$
has a unique expansion of the form
\begin{equation}
F=\sum_{\alpha \in \mathcal{J}}c_{\alpha }K_{\alpha }.
\label{10.14}
\end{equation}
with $c_{\alpha }\in \mathbb{R}$. 
Moreover,
\begin{equation}
\left\Vert F\right\Vert _{L^{2}(\P)}^{2}=\sum_{\alpha \in \mathcal{J}}\alpha
!c_{\alpha }^{2}.  \label{10.15}
\end{equation}
\end{theorem}

\bigskip
\begin{example}
\label{Ex10.3}
Let $h\in L^{2}(\mathbb{R})$ and define, for $x \in \R_+^n$,
\begin{align}
F=\int_0^{x} h(\xi)dL(\xi).
\end{align}
By  \eqref{LN} this can be written
\begin{align}
    F=I_{1}(h(\xi)\chi_{\Pi_{i=1}^n [0,x_i]}(\xi)\zeta).
\end{align}
Therefore $F$ has the following expansion, writing $\chi_{\Pi_{i=1}^n [0,x_i]}=\chi_{[0,x]}$ for short:
\begin{eqnarray}
F &=&I_{1}(\sum_{i\geq 1}(h\chi_{[0,x]},e _{i})e _{i}(\xi)\zeta)=\sum_{i\geq 1}(h\chi_{[0,x]},e
_{i})I_{1}(e _{i}(\xi)\zeta) \\
&=&\sum_{i\geq 1}\left( \int_0^x h(\xi)e _{i}(\xi)d\xi\right)
K_{\varepsilon ^{(i,1)}}m_{2}.  \nonumber
\end{eqnarray}
Recall that $z=m_{2}p_{1}(z);$ see (\ref{10.5}).
In particular, choosing $h=1$ we get the following expansion for the L\' evy sheet $L$:
\begin{equation}
L(x)=\sum_{i\geq 1}\left( \int_{0}^{x}e _{i}(\xi)d\xi\right)
K_{\varepsilon ^{(i,1)}}m_{2}.  \label{LK}
\end{equation}
\end{example}
In the following we put
\begin{equation}
(2\mathbb{N})^{k\alpha }=\prod_{j\geq 1}(2j)^{k\alpha _{j}},
\end{equation}%
if $\alpha =(\alpha _{1,}\alpha _{2,},...)\in \mathcal{J}$.

\begin{definition}
\label{Def10.4}{\bf (The L\'{e}vy-Hida spaces)} 

\begin{itemize}
\item[(i)] \quad\index{L\'evy-Hida stochastic test functin space}
{\bf Stochastic test functions $(\mathcal{S})$.} \newline
We define \ $(\mathcal{S})$ to be the set of all $%
\varphi =\sum_{\alpha \in \mathcal{J}}a_{\alpha }K_{\alpha }\in L^{2}(\P)$
such that%
\begin{equation}
\left\Vert \varphi \right\Vert _{k}^{2}:=\sum_{\alpha \in \mathcal{J}%
}a_{\alpha }^{2}\alpha !(2\mathbb{N})^{k\alpha }<\infty \text{ for \emph{all}
}k\in \mathbb{N},  \label{10.18}
\end{equation}%
equipped with the projective topology.%

\item[(ii)]\quad
{\bf Stochastic distributions $(\mathcal{S})^{* }$.} \newline
We define $(\mathcal{S})^{* }$ to be the set of all expansions $F=\sum_{\alpha \in \mathcal{%
J}}b_{\alpha }K_{\alpha }$ such that%
\begin{equation}
\left\Vert F\right\Vert _{-q}^{2}:=\sum_{\alpha \in \mathcal{J}}b_{\alpha
}^{2}\alpha !(2\mathbb{N})^{-q\alpha }<\infty \text{ for \emph{some} }q\in
\mathbb{N}.  \label{norm}
\end{equation}%
endowed with the inductive topology.\\

The space $(\mathcal{S})^{* }$ is the
dual of $(\mathcal{S}).$ If $F=\sum_{\alpha \in \mathcal{J}}b_{\alpha
}K_{\alpha }\in (\mathcal{S})^{* }$ and $\varphi =\sum_{\alpha \in
\mathcal{J}}a_{\alpha }K_{\alpha }\in (\mathcal{S}),$ then the action of $F$
on $\varphi $ is%
\begin{equation}
\left\langle F,\varphi \right\rangle =\sum_{\alpha \in \mathcal{J}}a_{\alpha
}b_{\alpha }\alpha !.  \label{10.21}
\end{equation}
\item[(iii)]\quad
{\bf Generalized expectation.}\newline
If  $F= \sum_{\alpha \in \mathcal{J}} a_{\alpha }K_{\alpha } \in (\mathcal{S})^{* }$, we define
the generalized expectation $\E[F]$ of $F$ by
$$
\E[F] = a_0.
$$ 
Note that $\E[K_{\alpha }] = 0$ for all $\alpha \ne 0$.
Therefore the generalized expectation coincides with the usual expectation if $F \in L^2(\P)$. 
\end{itemize}
\end{definition}

\bigskip\noindent

\subsection{White noise for L\' evy sheets and their compensated Poisson random measures}
Now we have the prerequisites to define the white noise of a L\'evy sheet and its associated Poisson random measure. In view of \eqref{LK} the following definition is natural:

\begin{definition} 
\begin{enumerate}
\item[(i)]
The \emph{L\'{e}vy white noise}
$$\overset{\bullet }{L
}(x):=V(x):= \frac{\partial^n L(x)}{\partial x_1 \partial x_2  ... ,\partial x_n}$$
is defined by the expansion
\begin{eqnarray*}
V(x):&=&m_{2}\sum_{i\geq 1}e _{i}(x)K_{\varepsilon
^{(i,1)}}=m_{2}\sum_{i\geq 1}e _{i}(x)I_{1}(e _{i}(x)p_{1}(z))\\
&=&m_{2}\sum_{i\geq 1}e _{i}(x)I_{1}(e _{i}(x)z).  
\end{eqnarray*}%
\item[(ii)]
The \emph{white noise} $\overset{\bullet }{\widetilde{N}%
}(x,z)$ of the compensated random measure  $\widetilde{N}(dx,dz)$ is defined by the expansion%
\begin{equation}
\overset{\bullet }{\widetilde{N}}(x,z)=\sum_{i,j\geq 1}e
_{i}(x)p_{j}(z)K_{_{\varepsilon ^{(i,j)}}}(\omega ).  
\end{equation}
\end{enumerate}
\end{definition}

\begin{remark}
\label{Rem10.6}Note that in the case of the expansion for $\overset{\bullet }{L}(x)$ we have%
\[
\sum_{\alpha \in \mathcal{J}}c_{\alpha }^{2}\alpha !(2\mathbb{N})^{-q\alpha
}=m_{2}^{2}\sum_{i\geq 1}e _{i}^{2}(x) 2^{-q}\kappa (i,1)^{-q}<\infty \text{ }
\]%
for $q\geq 2.$ Here we have used that $\kappa (i,1)=1+(i-1)i/2\geq i,$ and the following
well-known estimate for the Hermite functions:%
\begin{equation}
\sup_{x\in \mathbb{R}}\left\vert e _{m}(x)\right\vert =\mathcal{O}(m^{-\frac{1}{12}%
}).  
\end{equation}%
Therefore $\overset{\bullet }{L}(x)\in (\mathcal{S})^{* }$ for all $x.
$ Similarly $\overset{\bullet }{\widetilde{N}}(x,z)\in (\mathcal{S})^{* }$
for all $x,z.$
\end{remark}

\begin{remark}
\begin{enumerate}
\item[(i)]
By comparing the expansions%
\[
L(x)=\sum_{i\geq 1}\left( \int_{0}^{x_1}\int_0^{x_2} ... \int_0^{x_n} e _{i}(\xi) d\xi\right)
K_{\varepsilon ^{(i,1)}}m_{2}
\]%
and%
\[
\overset{\bullet }{L}(x)=m_{2}\sum_{i\geq 1}e _{i}(x)K_{\varepsilon
^{(i,1)}},
\]%
we get formally
\begin{equation}
\overset{\bullet }{L}(x)=\frac{\partial^n L(x)}{\partial x_1 \partial x_2 ..., \partial x_n}\text{ (derivative in }(%
\mathcal{S})^{* }). 
\end{equation}
This can be proved rigorously.
\item[(ii)] 
Choose a Borel set $\mho $ such that $\overline{\mho }\subset \mathbb{%
R}\backslash \{0\}.$ Then%
\begin{eqnarray*}
\widetilde{N}(x,\mho ) &=&I_{1}\left( \chi _{[0,x]}(\xi)\chi _{\mho
}(z)\right)  \\
&=&\sum_{i,j\geq 1}(\chi _{[0,x]},e _{i})_{L^{2}(\mathbb{R}^n)}(\chi
_{\mho },p_{j})_{_{L^{2}(\nu )}}I_{1}\left( e _{i}(\xi),p_{j}(\zeta)\right)  \\
&=&\sum_{i,j\geq 1}\int_{0}^{x}e _{i}(\xi)d\xi\cdot \int_{\mho }p_{j}(\zeta)\nu
(d\zeta)K_{\varepsilon ^{(i,j)}}(\omega ).
\end{eqnarray*}%
This justifies the identity
\begin{equation}
\overset{\bullet }{\widetilde{N}}(x,z)=\frac{\widetilde{N}(dx,dz)}{dx\times
\nu (dz)}\text{ (Radon-Nikodym derivative).}  \label{10.26}
\end{equation}%
\item[(iii)]
Also note that $\overset{\bullet }{L}$ is related to $%
\overset{\bullet }{\widetilde{N}}$ by%
\begin{equation}
\overset{\bullet }{L}(x)=\int_{\realio} \overset{\bullet }{\widetilde{N%
}}(x,\zeta)\nu (d\zeta).  
\end{equation}%
\dproof
To see this we consider%
\begin{eqnarray*}
&&\int_{\realio}\overset{\bullet }{\widetilde{N}}(x,\zeta)\zeta\nu (d\zeta) \\
&=&\int_{\realio}\sum_{i,j\geq 1}e _{i}(x)p_{j}(\zeta)K_{_{\varepsilon
^{(i,j)}}}(\omega )\zeta\nu (d\zeta)
=\sum_{i\geq 1}e _{i}(x)I_{1}( e
_{i}p_j)\sum_{j\geq 1}\int_{\realio} p_{j}(\zeta)\zeta\nu (d\zeta))  \\
&=&\sum_{i\geq 1}e _{i}(x)I_{1}\left( e _{i}(\xi)\zeta\right) =\overset{%
\bullet }{L}(t).
\end{eqnarray*}
\fproof
\end{enumerate}
\end{remark}

\section{Application: The stochastic heat equation with fractional time
driven by time-space Brownian and L\' evy white noise}
As an application of the multiparameter L\' evy white noise theory, we consider the fractional heat equation driven by time-space Brownian and L\' evy white noise. Our equation generalizes the fractional stochastic heat equation studied in \cite{MO1}, since here we account not only for the time-space Brownian noise but also for the L\' evy white noise. Furthermore, in the numerical example, we demonstrate that the L\' evy sheet plays an important role in modeling problems that are closer to real-life phenomena.
In this section, we are interested to the following fractional stochastic heat equation: 
\begin{equation}\label{heat1}
 	 \frac{\partial^{\alpha}}{\partial t^{\alpha}}Y(t,x)= \l \Delta Y(t,x)+\sigma W(t,x)+\gamma V(t,x);\; (t,x)\in (0,\infty)\times \mathbb{R}^{d}.
 	 	\end{equation}
       Here $d\in\mathbb{N}=\{1,2,...\}$ is the space dimension, $x \in \R^d$ denotes a point in space, $t\geq 0$ denotes the time,  $\frac{\partial^{\alpha}}{\partial t^{\alpha}}$ is the Caputo derivative of order $\alpha \in (0,2)$, and $\l>0$, $\sigma, \gamma \in \mathbb{R}$ are given constants,
   
  	\begin{equation}
  		\Delta Y =\sum_{j=1}^{d}\frac{\partial^{2}Y}{\partial x_{j}^{2}}(t,x)
  	\end{equation}
  is the Laplacian operator with respect to $x$, and
 
  \begin{equation}
  	W(t,x)=W(t,x,\omega)=\frac{\partial}{\partial t}\frac{\partial^{d}B(t,x)}{\partial x_{1}...\partial x_{d}} 
  \end{equation}
is time-space white noise, where $$B(t,x)=B(t,x,\omega); t\geq 0, x=(x_1, x_2, ... ,x_d) \in \R^d, \omega \in \Omega$$
 is time-space Brownian sheet with probability law $\P$, and

 \begin{align}
     V(t,x) = V(t,x,{\omega)=\frac{\partial}{\partial t}\frac{\partial^{d}L(t,x)}{\partial x_{1}...\partial x_{d}} }
 \end{align}
 is a L\'evy white noise, where $L(t,x)$ is a $1+d$ - parameter L\'evy sheet.
The boundary conditions are
\begin{align}
    Y(0,x)&=\delta(x)\text{ (the Dirac measure at  } x), \label{1.4}\\
    \lim_{x \rightarrow +/- \text{ }\infty}Y(t,x)&=0.\label{1.5}
\end{align}

\begin{remark}
    $W(t,x)$ and $V(t,x)$ model two different types of noise, which both can be represented as elements of $(\S)^{*}$: \\
    W (t, x) is a continuous Gaussian noise, while $V(t,x)$ represents noise coming from jumps from a L\' evy sheet. \\
    By construction we can assume that $W(t,x)$ and $V(t,x)$ are independent under the same probability measure $\P$ on $\Omega=\S'(\R_+ \times \R^d)$.
\end{remark}
By extending the method in \cite{MO1} to our equation, we obtain the following result, which we state without proof: 
\begin{theorem} \label{th1}
	The unique solution $Y(t,x) \in \mathcal{S}'$ of the fractional stochastic heat equation \eqref{heat1} - \eqref{1.5} is given by	
	\begin{align}\label{th2}
		Y(t,x)&=I_1 + I_2+I_3,
	\end{align}
	where
 		\begin{align}
			I_1=(2\pi)^{-d} \int_{\mathbb{R}^d} e^{ixy} E_{\alpha}(- \l t^{\alpha} |y|^2) dy
			=(2\pi)^{-d} \int_{\mathbb{R}^d} e^{ixy}\sum_{k=0}^{\infty} \frac{(- \l t^{\alpha} |y|^2)^k}{\Gamma(\alpha k +1)}dy, 
		\end{align}
	and
			\begin{align}
			I_2=\sigma (2\pi)^{-d} \int_{0}^{t}(t-r)^{\alpha -1}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}e^{i(x-z)y}\sum_{k=0}^{\infty}\frac{(-\l (t-r)^{\alpha}|y|^2)^{k}}{\Gamma(\alpha k+\alpha))}dy\right) B(dr,dz) , 
		\end{align}
        and
        \begin{align}
			I_3
			=\sigma (2\pi)^{-d} \int_{0}^{t}(t-r)^{\alpha -1}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}e^{i(x-z)y}\sum_{k=0}^{\infty}\frac{(-\l (t-r)^{\alpha}|y|^2)^{k}}{\Gamma(\alpha k+\alpha))}dy\right) L(dr,dz). 
		\end{align}
 
\end{theorem}

 \subsection{Application to Tumor Cell Invasion in Human Tissues}
 \subsection*{Introduction}

The dynamics of cancer invasion represent one of the complex and important problems in modern biology. It was stated by Murray in his seminal work on mathematical biology:

\begin{quote}
    ``The process of tumor invasion is not a simple, deterministic unfolding of a genetic program. It is a spatially structured, stochastic competition between the tumor and the host tissue, mediated by a myriad of biochemical and biomechanical signals. Capturing this requires models that can integrate deterministic drivers with intrinsic randomness and discrete, transformative events.'' \cite{murray2003}
\end{quote}

The important question is; what is the role of the multi-parameter L\'evy noise? A multi-parameter L\'evy sheet $V(t, x)$ allows us to model events that are localized in both space and time. A jump at a point $(t_0, x_0)$ represents a sudden, disruptive event happening at a precise location $x_0$ at the exact moment $t_0$. This is the strong point to simulating the spatially heterogeneous nature of real tumor.

\subsection*{Modeling the problem}

Our model for tumor cell density $Y(t, x)$ is:
\begin{equation}
\frac{\partial^\alpha}{\partial t^\alpha} Y(t, x) = \lambda \Delta Y(t, x) + \sigma W(t, x) + \gamma V(t, x)
\label{eq:tumor_complete_model}
\end{equation}

In this application, we model the invasion of a solid tumor through a host tissue, such as breast or brain parenchyma. \\
The solution $Y(t,x)$ represents the logarithm of the tumor cell density. The initial condition $Y(0,x)=\delta(x)$ represents a localized cluster of tumor cells initiating the invasion process.

\begin{itemize}
    \item \textbf{$\frac{\partial^\alpha}{\partial t^\alpha} Y$:} Models the ``anomalous diffusion'' of cells. We know that cells move in a complex random motion with memory path as a result of interactions with the tissue structure.\\
     - In the classical case, i. e. when $\alpha=1$, this equation models the normal diffusion of heat in a random  or noisy medium, the noise being represented by the time-space white noises $W(t,x)$ and $V(t,x)$. 

 - When $\alpha >1$ the inclusion can be used to model \emph{superdiffusion or enhanced diffusion}, where the particles spread faster than in regular diffusion. For example, this may occur in some biological systems.
 
 - When $\alpha <1$ the inclusion models \emph{subdiffusion}, in which the times of travel of the particles are longer than in the standard case. Such a situation may occur in transport systems. 
    
    \item \textbf{$\lambda \Delta Y$:} Models the random motility of cancer cells through the tissue. The coefficient $\lambda$ represents the cell motility rate, help to quantify how the active cells move and disperse from high.
    
    \item \textbf{$\sigma W(t, x)$:} Represents random fluctuations in the tumor environment, such as microscopic variations in nutrients, constantly drive the cells, making their behavior unpredictable.
    
    \item \textbf{$\gamma V(t, x)$:} models sudden events in the tumor microenvironment that can change the direction of invasion. See e.g.  \cite{fedotov2007}.
\end{itemize}

\end{document}